two cyclists who are 15 km apart start at the same time traveling the same direction. after 6 hours the faster one overtakes the slower one. if they had traveled toward each other, they would have met after 2 hrs. What is the rate of each cyclist?
Answer by TimothyLamb(4379)

x = speed of faster biker
y = speed of slower biker
---
s = d/t
---
relative speed of the bikers traveling in the same direction:
x - y = 15/6
x - y = 2.5 kph
---
relative speed of the bikers traveling towards each other:
x + y = 15/2
x + y = 7.5 kph
---
put the system of linear equations into standard form
---
x - y = 2.5
x + y = 7.5
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = speed of faster biker = 5 kph
y = speed of slower biker = 2.5 kph
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---